. The force R is the resultant of the forces P1, P2, and P3 acting on the rectangular
plate. Find P1 and P2 if
1
Expert's answer
2021-10-11T07:01:25-0400

"x-component \\
R_x=\sum F_x\\
40 \cos 30= \frac{1}{\sqrt{5}}P_1+\frac{3}{5}P_2-20\\
54.641= \frac{1}{\sqrt{5}}P_1+\frac{3}{5}P_2...1\\

y-component \\
R_y=\sum F_y\\
40 \sin 30= \frac{2}{\sqrt{5}}P_1+\frac{4}{5}P_2\\
20= \frac{2}{\sqrt{5}}P_1+\frac{4}{5}P_2...2\\

P_1=62.3 kN\\

P_2=44.6 kN\\"